OpenStudy (anonymous):

Five friends agree to save money for a graduation road trip. They decide that each of them will put $0.25 in the fund on the first day of May, $0.50 on the second day, $0.75 on the third day, and so on. At the end of May, there will be $_____ in their fund. (Hint: There are 31 days in May.)

OpenStudy (anonymous):

arithmetic series S(n)=(n/2)*(a1+a(n)) 1.25, 2.50, 3.75... S(31)=(31/2)(1.25+a(n)) a(n)=1.25+(31-1)*1.25=38.75 S(31)=15.5*(1.25+38.75)=620